OpenStudy (anonymous):

using substitution or elimination solve this: x^2-y^2=15 and x+y=1

OpenStudy (anonymous):

Notice that x^2 - y^2 is a difference of two squares. This means: x^2-y^2 = (x+y)*(x-y) = (1)*(x-y) = x-y = 15. Therefore you have two equations and two unknowns: 1) x-y=15 2) x+y=1

OpenStudy (anonymous):

can it still be solved?

OpenStudy (anonymous):

can you continue the problem and use elimination?

OpenStudy (anonymous):

or is this no solution?